# Adaptive Testing for LLM Evaluation: A Psychometric Alternative to Static Benchmarks

### v1

@@ Abstract
Large language model evaluation requires thousands of benchmark items, making evaluations expensive and slow. Existing methods compute average accuracy across fixed item sets, treating all items equally despite varying quality and informativeness. We present ATLAS  an adaptive testing framework using Item Response Theory (IRT) to estimate model ability through Fisher information-guided item selection. Our analysis of five major benchmarks reveals that 3-6% of items exhibit negative discrimination, indicating annotation errors that corrupt static evaluation. ATLAS achieves 90% item reduction while maintaining measurement precision: on HellaSwag (5,608 items), we match full-benchmark estimates using only 42 items with 0.154 MAE. Our framework maintains item exposure rates below 10% and test overlap at 16-27%, compared to static benchmarks where every model sees all items (100% exposure). Among 4,000+ tested models, IRT ranks differ from accuracy ranks: models with the same accuracy get different IRT scores, and 23-31% of all models shift by more than 10 rank positions. Code and calibrated item banks are available at https://github.com/Peiyu-Georgia-Li/ATLAS.git.
@@ 1 Introduction
Large language model (LLM) evaluation relies on benchmarks with tens of thousands of test items. These benchmarks impose high computational costs and evaluation cycles spanning days or weeks. Despite benchmark growth—some proposals exceed 100,000 items—evaluation practice remains static: models receive average accuracy scores across fixed item sets. This approach ignores statistical information in large datasets and raises questions about efficiency and validity.
@@
Current evaluation faces three fundamental limitations. First, average scores hide meaningful differences between models with distinct error patterns, particularly for lower-performing models where small ability differences are obscured by measurement noise. Second, static benchmarks create vulnerability to data contamination as items leak into pretraining corpora, enabling high scores through memorization rather than genuine capability. Third, evaluating complete benchmarks is inefficient, treating poorly discriminative items as equally informative as high-quality questions.
@@
To address these limitations, we propose ATLAS (Adaptive Testing for LLM Ability Scoring), an adaptive evaluation framework based on computerized adaptive testing (CAT; (lord1980, ; wainer2000computerized, ; weiss1982improving, )). ATLAS first calibrates benchmark items using three-parameter logistic (3PL) IRT models to estimate item difficulty, discrimination, and guessing parameters (birnbaum1968some, ; hambleton1991fundamentals, ). Then, rather than administering fixed item sets, ATLAS dynamically selects items with maximum Fisher information for each model’s current estimated ability, terminating when precision thresholds are reached. This approach directly addresses all three limitations: Fisher information-guided selection provides precise ability estimates that distinguish models with identical accuracy, dynamic item selection reduces contamination through diverse test forms, and adaptive termination achieves reliable evaluation with significantly fewer items.
We validate our framework across five major benchmarks (WinoGrande, TruthfulQA, HellaSwag, GSM8K, ARC). ATLAS achieves competitive or superior performance to static baselines: on TruthfulQA, it attains the lowest MAE (0.067) using only 51 items; on HellaSwag, it matches MetaBench accuracy (0.154 MAE) while requiring 3× fewer items (39 vs 93). Across benchmarks, ATLAS requires 30-78 items on average compared to thousands in full evaluation, while IRT-based estimates reveal systematic rank reordering that exposes limitations of accuracy-based evaluation.
Our contributions are: (1) We demonstrate fundamental limitations of average-score evaluation and establish the benefits of psychometric ability estimates for LLM evaluation. (2) We develop ATLAS, the first large-scale adaptive testing framework for LLMs, achieving 90% item reduction while maintaining measurement precision and natural contamination resistance through $10\%$ item exposure rates. (3) We provide comprehensive psychometric analysis of five major benchmarks, revealing that IRT-based estimates cause systematic rank reordering ($23-31\%$ of models shift more than 10 positions) and identifying that $3-6\%$ of items exhibit negative discrimination due to annotation errors, with calibrated item banks released for reproducible evaluation.
@@ 2 Related Work
2.1 Background
Large-scale benchmarks such as WinoGrande, TruthfulQA, HellaSwag, GSM8K and ARC have become standard for LLM evaluation but impose significant computational costs. These static approaches suffer from benchmark saturation, data contamination, and item redundancy, raising concerns about their ability to reflect genuine model capabilities.
Item Response Theory (IRT) has recently been applied to LLM evaluation (lalor2024item, ; guinet2025evaluating, ). It provides item parameters such as difficulty, discrimination, and guessing, as well as latent ability estimates $\theta$ for models. However, existing IRT applications remain largely static: tinyBenchmarks (polo2024tinybenchmarks, ) uses clustering for item selection but doesn’t guarantee informativeness for $\theta$ estimation, while MetaBench (metabench, ) requires computationally expensive iterations to identify stable subsets. These approaches often lack proper psychometric validation and emphasize predictive accuracy over model fit diagnostics, which makes it difficult to ensure that ability estimates are valid, interpretable, and comparable across models.
@@ 2.2 Adaptive Testing
Computerized adaptive testing (CAT) is a psychometric framework for efficiently measuring ability by selecting items that maximize Fisher information. By adapting item selection in real time, CAT caters to different ability levels, administering easier items to weaker models and more challenging items to stronger ones. CAT can operate with a fixed test length or with a variable length, terminating once a predefined precision threshold is reached (weiss1982improving, ).
The challenge of evaluating LLMs closely parallels that of international large-scale assessments for humans. Our dataset includes more than 4,000 models, ranging from simple pattern-matching systems with very low ability estimates ($\theta<-2$) to advanced models such as GPT-4 with high ability estimates ($\theta>2$) . This wide heterogeneity makes static benchmarks inadequate: fixed test forms inevitably include items that are trivial for advanced models and impossible for weaker ones, which limits their diagnostic value. Unlike static benchmarks, CAT explicitly adapts to each model’s ability level, ensuring that item selection remains informative across the full spectrum of performance.
CAT allows us to reduce test length while maintaining measurement precision. Modern extensions such as Multi-Stage Testing and the incorporation of process data (zenisky2009multistage, ; zheng2015fly, ; tang2024utilizing, ) provide further gains in efficiency and validity. Prior attempts to explore adaptive evaluation for LLMs have been either limited in scope (zhuang2023efficiently, ) or primarily conceptual (zhuangposition, ). Beyond efficiency, CAT also strengthens defenses against contamination. Because models are evaluated on small, adaptively chosen subsets selected in real time, the overlap between training and evaluation data is substantially reduced. Furthermore, CAT protocols can be refreshed with new items while preserving comparability of ability estimates $\theta$. A detailed comparison of IRT-based approaches is provided in Appendix B.
@@ 3 Methodology
We introduce a novel adaptive testing framework that transforms LLM evaluation from static benchmarking to dynamic ability estimation. Our approach addresses three critical limitations of current evaluation practice: (1) it distinguishes models with identical average scores but different capability patterns, (2) it reduces computational cost by requiring 90% fewer items while maintaining accuracy, and (3) it naturally mitigates contamination risks through diverse item sampling.
@@ 3.1 Problem Formulation and Setup
3.2 Discriminative Item Filtering
@@
Model Selection. We retain only models $\mathcal{L}$ with complete responses across all items to ensure unbiased ability estimates. Models with extreme scores (below 0.1st percentile) are excluded to prevent parameter estimation instability, as IRT’s sigmoidal functions become under-constrained at the boundaries.
Item Filtering. We apply two complementary filters to retain only discriminative items:
  • Low-variance removal: Items with response standard deviation $1\%$ or mean accuracy $95\%$ are discarded, as they provide litter information for differentiating between models.
  • Discrimination filtering: We compute the point-biserial correlation $r_{pb}(i)$ between each item’s response vector $\{Y_{i,\ell}\}_{\ell\in\mathcal{L}}$ and the models’ total scores $T_{\ell}=\sum_{j\in\mathcal{I}}Y_{j,\ell}$ (see Appendix E for details). Items with $r_{pb}(i)<0.1$ are removed as non-diagnostic.
This filtering process yields a refined response matrix that supports stable and reliable IRT calibration (see Table 5 in Appendix E for detailed results).
@@ 3.3 Scalable IRT Calibration
The calibration stage estimates item parameters $(a_{i},b_{i},c_{i})$ and computes reference ability estimates $\hat{\theta}_{\ell}^{\mathrm{whole}}$ for each LLM $\ell$ for validation. To model the probability of a correct response, we adopt the three-parameter logistic (3PL) IRT model (birnbaum1968some, ; lord1980, ):
@@
Common-Person Calibration at Scale. We develop a partition-based calibration strategy that takes advantage of the unique characteristics of LLM benchmarking. Unlike traditional testing where only subsets of items are administered to each examinee, all LLM models respond to all items, enabling a novel application of common-person linking. We partition items into $K$ non-overlapping subsets $\mathcal{I}_{k}$ (each with $|\mathcal{I}_{k}|\geq 100$ items), calibrate each subset independently, then apply mean-sigma transformations (kolen2014, ) using the model population $\mathcal{L}$ as linking anchors. This approach reduces computational complexity from $O(|\mathcal{I}|^{3})$ to $O(K\cdot\max_{k}|\mathcal{I}_{k}|^{3})$ while maintaining calibration accuracy through the redundancy of having all models as linkers (chalmers2012mirt, ).
Heterogeneity-Aware Ability Estimation. The extreme heterogeneity of LLM populations, spanning from random baselines ($\theta\approx-3$) to near-perfect models ($\theta\approx 3$), breaks standard estimation methods. Standard IRT practice relies on Expected A Posteriori (EAP) estimation (bock1982adaptive, ) for ability estimation. However, EAP produces undefined values for all-correct or all-incorrect response patterns, affecting 12% of our model population spanning from random baselines to near-perfect systems. To address this LLM-specific challenge, we employ Weighted Likelihood Estimation (WLE) (warm1989weighted, ) with bias correction term $\frac{J(\theta)}{2I(\theta)}$ where $J(\theta)=\sum_{i}\frac{\partial I_{i}(\theta)}{\partial\theta}$ (warm1989weighted, ). This choice enables finite estimates across the full ability spectrum while maintaining consistency properties (baker2004item, ), crucial for establishing reliable validation baselines in the extreme heterogeneity of LLM evaluation.
Multi-Subset Model Fit Validation. Unlike prior IRT applications to LLM evaluation (polo2024tinybenchmarks, ; metabench, ) that omit model validation, we conduct rigorous psychometric diagnostics to ensure calibration quality. We report the $M_{2}$ statistic (maydeu2015m2, ) with RMSEA indices, establishing that our 3PL models achieve good fit (RMSEA <0.05<0.05) across all benchmarks, as summarized in Appendix D. Our partition-based calibration strategy, which divides the full item bank into non-overlapping subsets (each containing $\geq 100$ items for statistical stability), enables both computational feasibility and robust validation. Since the same set of models $\mathcal{L}$ acts as common persons across all partitions, diagnostic statistics reflect global calibration quality rather than partition-specific artifacts. This multi-subset linking design ensures that model fit metrics capture systematic patterns across the entire item bank, not just localized subsets. This validation is crucial for reliable adaptive testing, as misfitting items would compromise Fisher information calculations and degrade selection accuracy.
@@ 3.4 Adaptive Testing with Information Selection
Initialization and Bounds. We initialize the ability estimate at $\hat{\theta}_{0}=0$. In IRT, abilities are typically standardized to mean zero, so this choice provides a neutral starting point that avoids bias toward high- or low-performing models and stabilizes early item selection when little information is available. We enforce minimum (30) and maximum (500) item limits. The minimum ensures stable estimation for models at performance extremes, while the maximum constrains computational cost and yields approximately 90% reduction in test length relative to full benchmarks.
Randomesque Item Selection. Rather than deterministically selecting the single most informative item, we randomly sample from the top-5 candidates ranked by Fisher information:
$I_{i}(\theta)=a_{i}^{2}\cdot p_{i}(\theta)\cdot[1-p_{i}(\theta)].$ (2)
This randomesque strategy (kingsbury1989procedures, ) prevents over-reliance on specific item types while still keeping high information, which is important for models with specialized capabilities.
Sequential Ability Updates. After each item administration, we update the ability estimate using Expected A Posteriori (EAP) estimation (bock1982adaptive, ):
@@
Output and Validation. For each model $\ell$, the algorithm produces: (1) the administered item sequence and responses $R_{\ell}$, (2) the ability estimate trajectory $\{\hat{\theta}_{t}\}$ with associated standard errors, and (3) the final estimate $\hat{\theta}_{\ell}$. We validate these adaptive estimates against whole-bank references $\hat{\theta}_{\ell}^{\mathrm{whole}}$ to confirm that our dramatic reduction in items does not compromise measurement accuracy.

### v2

@@ Abstract
Evaluating large language models (LLMs) typically requires thousands of benchmark items, making the process expensive, slow, and increasingly impractical at scale. Existing evaluation protocols rely on average accuracy over fixed item sets, treating all items as equally informative despite substantial variation in difficulty and discrimination. We introduce ATLAS, an adaptive testing framework based on Item Response Theory (IRT) that estimates model ability using Fisher information–guided item selection. ATLAS reduces the number of required items by up to 90% while maintaining measurement precision. For instance, it matches whole-bank ability estimates using only 41 items (0.157 MAE) on HellaSwag (5,600 items). We further reconstruct accuracy from ATLAS’s ability estimates and find that reconstructed accuracies closely match raw accuracies across all five benchmarks, indicating that ability $\theta$ preserves the global performance structure. At the same time, $\theta$ provides finer discrimination within accuracy-equivalent models: among more than 3,000 evaluated models, 23–31% shift by more than 10 rank positions, and models with identical accuracies receive meaningfully different ability estimates. Code and calibrated item banks available at https://github.com/Peiyu-Georgia-Li/ATLAS.git.
@@ 1 Introduction
Large language model evaluation relies on benchmarks with tens of thousands of items, which are costly to run and often take days or weeks to complete. Even with benchmarks exceeding 100,000 items, evaluation still depends on average accuracy over fixed item sets. This practice overlooks valuable statistical information and raises concerns about efficiency and validity.
@@
Current evaluation practices face three fundamental limitations. First, average accuracy over fixed item sets obscure meaningful differences between models with distinct error patterns, especially among lower-performing models where small ability differences are dominated by measurement noise. Second, static evaluations treat poorly discriminative items as equally informative as high-quality questions, leading to unreliable and often misleading comparisons. Third, evaluating complete benchmarks is inefficient and time-consuming, requiring models to answer hundreds or thousands of items regardless of how much additional information those items provide.
@@
To address these limitations, we propose ATLAS (Adaptive Testing for LLM Ability Scoring), an adaptive evaluation framework based on computerized adaptive testing (CAT) (Lord, 2012; Wainer et al., 2000; Weiss, 1982). ATLAS first calibrates benchmark items using three-parameter logistic (3PL) IRT models to estimate item difficulty, discrimination, and guessing parameters (Birnbaum, 1968; Hambleton et al., 1991). Then, rather than administering fixed item sets, ATLAS dynamically selects items with maximum Fisher information for each model’s current estimated ability, terminating when precision thresholds are reached. This approach directly addresses all three limitations: Fisher information-guided selection provides precise ability estimates that distinguish models with identical accuracy, dynamic item selection prioritizes highly discriminative items rather than treating all questions as equally informative, and adaptive termination enables reliable evaluation with far fewer items and substantially less time than full-benchmark scoring.
We evaluate ATLAS across five major benchmarks, including WinoGrande (Sakaguchi et al., 2021), TruthfulQA (Lin et al., 2022), HellaSwag (Zellers et al., 2019), GSM8K (Cobbe et al., 2021), ARC (Clark et al., 2018) and find that it matches or exceeds the accuracy of strong static baselines while using far fewer items. For example, ATLAS achieves the lowest MAE on TruthfulQA ($0.064$ with 48 items) and HellaSwag ($0.157$ with 41 items), and matches MetaBench (Kipnis et al., 2025) on WinoGrande while using 2× fewer items (70 vs. 133). It also outperforms TinyBenchmarks (Polo et al., 2024), which uses 97–100 items but yields higher error across all benchmarks. Overall, ATLAS requires only 30–89 items per benchmark compared to hundreds in static subsets, and attains the lowest Information Efficiency Score (IES) across all benchmarks, demonstrating the strongest accuracy–efficiency tradeoff.
Our contributions are: (1) We identify fundamental limitations of average-score evaluation and show that psychometric ability estimates provide more robust and informative comparisons of LLM performance. (2) We introduce ATLAS, a large-scale adaptive testing framework for LLMs that achieves up to 90% item reduction while maintaining measurement precision through SE-controlled stopping, enabling flexible and precision-targeted evaluation beyond fixed-length designs. (3) We conduct a comprehensive psychometric analysis of five major benchmarks, revealing that IRT-based ability estimation induces substantial rank reordering (23–31% of models shift by more than 10 positions). (4) We highlight the importance of rigorous psychometric validation by reporting model-fit statistics (e.g., RMSEA via the M2 statistic) and demonstrating the use of common-person linking to align item parameters efficiently and ensure cross-model comparability.
@@ 2 Related Work
2.1 IRT-based Approaches
Item Response Theory (IRT) has recently been applied to LLM evaluation (Lalor et al., 2024). It provides item parameters such as difficulty, discrimination, and guessing, as well as latent ability estimates $\theta$ for models. However, existing IRT applications remain largely static in nature. For instance, TinyBenchmarks (Polo et al., 2024) uses clustering for item selection but doesn’t guarantee informativeness for $\theta$ estimation, while MetaBench (Kipnis et al., 2025) requires computationally expensive iterations to identify stable subsets. Moreover, these approaches often lack proper psychometric validation and emphasize predictive accuracy over model fit. TinyBenchmarks and MetaBench do not report fit statistics. Instead, we compute these metrics using their released IRT code (as shown in Table 1). This limitation makes it difficult to ensure that the resulting ability estimates are valid, interpretable, and comparable across models. A detailed comparison of IRT-based approaches is provided in Appendix B.
@@ 2.2 Adaptive Testing
Computerized adaptive testing (CAT) adjusts item administration based on an examinee’s evolving ability estimate (Meijer and Nering, 1999; Van der Linden and Glas, 2010). After each response, the test updates the ability estimate and selects the next item using an algorithm that aims to provide the most informative measurement while satisfying test constraints (Weiss, 1982; Chang, 2015; Cheng and Chang, 2009). Related adaptive frameworks such as multistage testing and process-data–based approaches apply similar principles and offer additional flexibility and diagnostic information (Zenisky et al., 2009; Zheng and Chang, 2015; Tang et al., 2025). These features allow CAT to evaluate examinees efficiently while maintaining rigorous measurement precision. This adaptive structure aligns well with challenges in evaluating LLMs, which vary widely in their performance levels. Current evaluations often use large static benchmarks in which every model must answer all items, even when many items provide little information about its ability. These benchmarks also rarely report empirical item characteristics, so their difficulty range and informativeness across models remain unclear. CAT addresses these limitations by selecting items targeted to each model’s estimated ability, which yields more precise and efficient evaluation with far fewer items.
However, only a few studies have explored adaptive evaluation for LLMs. Early efforts were either limited in scope (Zhuang et al., 2023) or primarily conceptual (Zhuang et al., 2025). A recent study that is closely related to our work is Fluid Benchmarking (Hofmann et al., 2025), which appeared around the same time as this research. Fluid also applies CAT principles to increase evaluation efficiency and models LLM performance on a latent ability scale. The two approaches are complementary and differ in several ways. Fluid focuses on adaptive evaluation during LLM pretraining, whereas our work examines post-training evaluation. Fluid calibrates its IRT model on 102 LMs, while ATLAS uses a substantially larger and more diverse pool of 3,000+ LMs. Fluid adopts a fixed-length adaptive design, while ATLAS uses a precision-based stopping rule that terminates the test once the uncertainty of the ability estimate falls below a predefined threshold. Precision-based stopping ensures consistent measurement precision across models and avoids administering unnecessary items. In addition, we report model-fit statistics to ensure the adequacy of the IRT model before running CAT and provide a transparent description of calibration and linking procedures used to estimate item parameters. Both studies demonstrate that adaptive testing methods can be used at different stages of LLM development and under different design choices, which illustrates the broader potential of CAT-based approaches for scalable and precise LLM assessment.
@@ 3 Methodology
We introduce a novel adaptive testing framework that transforms LLM evaluation from static benchmarking to dynamic ability estimation. Our approach addresses three critical limitations of current evaluation practice: (1) it reduces computational cost by requiring 90% fewer items while maintaining overall performance, (2) it overcomes the ceiling effects of accuracy-based metrics and preserves discrimination across the ability spectrum, and (3) it distinguishes models with identical accuracy score but different abilities.
@@ 3.1 Problem Formulation and Setup
3.2 Data Construction with psychometric filtering
@@
Model Selection and Splitting.
We retain only models $\mathcal{L}$ with complete responses across all items. To obtain a calibration sample whose ability distribution approximates a Gaussian, a standard assumption for stable IRT estimation, we exclude models in the extreme low-ability tail (below the 0.1st percentile), whose near-zero response patterns destabilize 3PL parameter estimation. The high-ability tail is small and non-degenerate, so these models are retained. The selected models are then split into training and testing sets using stratified random sampling (10 bins) to ensure that both splits share a similar ability distribution. We allocate 90% of the models to the training set for item calibration, and use the remaining 10% as the testing set for evaluating performance in our experiments (see Table 5).
Item Filtering.
We apply two complementary filters to retain only discriminative items:
  • Low-variance removal: Items with response standard deviation $1\%$ or mean accuracy $95\%$ are discarded, as they provide little information for differentiating between models.
  • Discrimination filtering: We compute the point-biserial correlation $r_{pb}(i)$ between each item’s response vector $\{Y_{i,\ell}\}_{\ell\in\mathcal{L}}$ and the models’ total scores $T_{\ell}=\sum_{j\in\mathcal{I}}Y_{j,\ell}$ (see Appendix C for details). Items with $r_{pb}(i)<0.1$ are removed as non-diagnostic.
This filtering process yields a refined response matrix that supports stable and reliable IRT calibration (see Table 5 for detailed results).
@@ 3.3 Scalable IRT Calibration
The calibration stage estimates item parameters $(a_{i},b_{i},c_{i})$ and computes reference ability estimates $\hat{\theta}_{\ell}^{\mathrm{whole}}$ for each LLM $\ell$ for validation. To model the probability of a correct response, we adopt the three-parameter logistic (3PL) IRT model (Birnbaum, 1968; Lord, 2012):
@@
Common-Person Calibration at Scale.
To estimate item characteristics efficiently using the 3PL model, we adopted a partition-based calibration procedure that leverages the unique structure of LLM benchmarking. Instead of fitting the full 3PL model to the entire item pool at once, which would be computationally prohibitive, we divided the items into $K$ non-overlapping subsets $\mathcal{I}_{k}$ (each with $|\mathcal{I}_{k}|\geq 100$ items), and calibrated each subset independently. This yields multiple provisional difficulty scales that must be aligned. Because all models answer all items, the model population serves as a natural set of common persons, allowing us to link the independently calibrated subsets onto a unified scale using mean–sigma transformations (Brennan, 2014). This approach reduces computational complexity from $O(|\mathcal{I}|^{3})$ to $O(K\cdot\max_{k}|\mathcal{I}_{k}|^{3})$ while maintaining calibration accuracy due to the stability provided by having all models serve as linking anchors (Chalmers, 2012). A detailed description is provided in Appendix C.2.
Heterogeneity-Aware Ability Estimation.
LLM populations exhibit extreme heterogeneity, ranging from near-random models ($\theta\approx-3$) to highly capable systems ($\theta\approx 3$). To obtain stable and unbiased estimates across this wide ability spectrum, we adopt the Weighted Likelihood Estimator (WLE) (Warm, 1989), which incorporates a bias-correction term $\frac{J(\theta)}{2I(\theta)}$, where $J(\theta)=\sum_{i}\frac{\partial I_{i}(\theta)}{\partial\theta}$. WLE provides finite, well-behaved estimates even at ability extremes and maintains desirable consistency properties (Baker and Kim, 2004). These characteristics are essential for establishing reliable evaluation baselines under the substantial heterogeneity present in modern LLM benchmarks.
Multi-Subset Model Fit Validation.
Unlike prior IRT-based LLM benchmarks (Polo et al., 2024; Kipnis et al., 2025), which do not report model-fit diagnostics, we perform explicit psychometric validation to ensure calibration quality. We compute the limited-information $M_{2}$ statistic with RMSEA (Maydeu-Olivares, 2013) for TinyBenchmarks, MetaBench, and ATLAS (Table 1). TinyBenchmarks exhibits severe misfit across all datasets, consistent with its underidentified calibration setting—estimating up to 15 latent traits from only 395 models. MetaBench performs better (RMSEA 0.04–0.14) but still shows poor fit on TruthfulQA and marginal fit on ARC, indicating uneven generalization across domains. In contrast, ATLAS consistently achieves acceptable or good fit across all benchmarks, demonstrating stable and robust calibration. Our partition-based calibration further supports reliable validation by dividing the item bank into non-overlapping subsets (each $\geq$100 items) while using a shared model set as common persons. This linking strategy ensures that fit statistics reflect global calibration quality rather than subset-specific artifacts, which is essential for accurate Fisher-information–based adaptive testing.
Method Winogrande TruthfulQA HellaSwag GSM8K ARC
RMSEA Fit RMSEA Fit RMSEA Fit RMSEA Fit RMSEA Fit
TinyBenchmarks 364.24 Poor 371.49 Poor 646.82 Poor 506.60 Poor 369.89 Poor
MetaBench 0.0524 Accept. 0.1389 Poor 0.0498 Good 0.0423 Good 0.0811 Marginal
ATLAS 0.0565 Accept. 0.0690 Accept. 0.0482 Good 0.0438 Good 0.0595 Accept.
Table 1: Model fit comparison across benchmarks using the limited-information statistic $M_{2}$ and its derived Avg. RMSEA values. Lower Avg. RMSEA indicates better model fit. Model fit is interpreted according to standard psychometric thresholds: RMSEA << 0.05 = Good fit; 0.05–0.08 = Acceptable fit (abbreviated as Accept.); 0.08–0.10 = Marginal fit; >> 0.10 = Poor fit.
@@ 3.4 Adaptive Testing with Information Selection
Initialization and Bounds.
We initialize the ability estimate at $\hat{\theta}_{0}=0$. This value is a conventional neutral starting point in CAT because the latent trait scale is typically assumed to be centered at zero. Beginning at the scale midpoint helps stabilize early item selection by preventing the algorithm from drifting toward artificially high or low values before any response information is available. We enforce minimum (30) and maximum (500) item limits. The minimum ensures stable estimation for models at performance extremes, while the maximum constrains computational cost and yields approximately 90% reduction in test length relative to full benchmarks.
Randomesque Item Selection.
Rather than deterministically selecting the single most informative item, we randomly sample from the top-5 candidates ranked by Fisher information:
$I_{i}(\theta)=\frac{\big(p_{i}^{\prime}(\theta)\big)^{2}}{p_{i}(\theta)\,\big(1-p_{i}(\theta)\big)}$ (2)
This randomesque strategy (Kingsbury and Zara, 1989) prevents over-reliance on specific item types while still keeping high information, which is important for models with specialized capabilities.
Sequential Ability Updates.
After each item administration, we update the ability estimate using Expected A Posteriori (EAP) estimation (Bock and Mislevy, 1982):
@@
Output and Validation.
For each model $\ell$, the algorithm produces: (1) the administered item sequence and responses $R_{\ell}$, (2) the ability estimate trajectory $\{\hat{\theta}_{t}\}$ with associated standard errors, and (3) the final estimate $\hat{\theta}_{\ell}$. We validate these adaptive estimates against whole-bank references $\hat{\theta}_{\ell}^{\mathrm{whole}}$ to confirm that our dramatic reduction in items does not compromise measurement accuracy.
(a) ARC
(b) TruthfulQA
(c) HellaSwag
(d) GSM8K
(e) WinoGrande
Figure 1: Comparison of subset (predicted) ability estimates against whole-bank (reference) abilities across five benchmarks (graphical illustration complementing Table 2). Points along the identity line indicate perfect agreement. ATLAS maintains the closest alignment overall, particularly on TruthfulQA, ARC and HellaSwag and in the high-ability regime of WinoGrande, while static baselines such as TinyBenchmarks and Metabench show greater variance and systematic deviation.
Method WinoGrande TruthfulQA HellaSwag
[…]
Method WinoGrande TruthfulQA HellaSwag
[…]